I am painting the four walls and a ceiling of a room. The room is 600cm long by 450cm wide and 300 cm high. What is the area?
Answer by Alan3354(69443)

I am painting the four walls and a ceiling of a room. The room is 600cm long by 450cm wide and 300 cm high. What is the area?
----------
2 walls are 6 m by 3 m = 18*2 = 36 sq meters
2 walls are 4.5 by 3 m = 13.5*2 = 27 sq meters
The ceiling is 6 m by 4.5 m = 27 sq meters
---------
Add those.